### v2

# Investigating Length Issues in Document-level Machine Translation

Abstract
Transformer architectures are increasingly effective at processing and generating very long chunks of texts, opening new perspectives for document-level machine translation (MT). In this work, we challenge the ability of MT systems to handle texts comprising up to several thousands of tokens. We design and implement a new approach designed to precisely measure the effect of length increments on MT outputs. Our experiments with two representative architectures unambiguously show that (a) translation performance decreases with the length of the input text; (b) the position of sentences within the document matters, and translation quality is higher for sentences occurring earlier in a document. We further show that manipulating the distribution of document lengths and of positional embeddings only marginally mitigates such problems. Our results suggest that even though document-level MT is computationally feasible, it does not yet match the performance of sentence-based MT.
Investigating Length Issues in Document-level Machine Translation
Ziqian Peng1,2, Rachel Bawden2, François Yvon1, 1Sorbonne Université & CNRS, ISIR, Paris, France, 2Inria, Paris, France, {ziqian.peng,francois.yvon}@isir.upmc.fr     rachel.bawden@inria.fr
footnotetext: ©  2025 The authors. This article is licensed under a Creative Commons 4.0 licence, no derivative works, attribution, CC-BY-ND.
1 Introduction
Statistical and neural machine translation (MT) architectures (Koehn, 2020) have been designed to process isolated sentences, limiting their ability to properly handle discourse phenomena, such as coherence and cohesion, the modelling of which requires longer contexts (Fernandes et al., 2023). A first step to address this shortcoming has been to augment the source and/or the target side with a couple of preceding sentences (Tiedemann and Scherrer, 2017). Multiple approaches to encode and fully exploit such extended contexts have been proposed (Popescu-Belis, 2019; Maruf et al., 2021; Castilho and Knowles, 2024)disabledisabletodo: disablecite more, the survey in 2024? and have been shown to improve the ability of MT engines to preserve local discourse coherence and cohesiveness through word-sense disambiguation or the resolution of anaphoric references (Bawden et al., 2018; Voita et al., 2018). Most of these approaches continue to process texts on a per-sentence basis with an extended context, even though attempts have also been made to process continuous chunks of texts comprising several sentences (Scherrer et al., 2019; Lopes et al., 2020; Ma et al., 2020, 2021; Lupo et al., 2022a; Wu et al., 2023).disabledisabletodo: disablecite more?
The ability of today’s neural MT models—relying on encoder-decoder or decoder-only architectures—to handle large context lengths, up to thousands of tokens (Peng et al., 2024), opens new perspectives to go beyond context-augmented MT and develop fully-fledged document-level MT, where the entire document context is available at once, and where the target text is generated in a single pass. [^1] These perspectives are, for instance, explored in the latest edition of the WMT shared task on General Machine Translation (Kocmi et al., 2024). Two main technical novelties have made this possible: (a) more efficient computation in the attention layers (Tay et al., 2022) and (b) changes in the design of positional encodings (PEs). In particular, replacing the sinusoidal absolute PEs (APEs) of (Vaswani et al., 2017) with methods like ALIBI Press et al. (2022) and RoPE (Su et al., 2024), which lend themselves well to length extrapolation (Sun et al., 2023; Zhao et al., 2024), seems to make today’s transformers amenable to the processing of arbitrarily long contexts (Mohtashami and Jaggi, 2023; Han et al., 2024).
In this work, we challenge the ability of contemporary MT models to effectively handle long spans of texts. For this, we develop a new methodology for assessing the impact of length variations on MT performance. We perform a series of controlled experiments with two representative neural MT systems, where the same documents are processed by chunks of increasing lengths in a document-level manner and show that (a) MT performance tends to degrade with the length of the source document, (b) length issues happen even for in-distribution lengths and get worse when extrapolating to unseen document lengths, and (c) most of the degradation happens in the final parts of the translation. Hypothesising that this may be due to a mismatch between the distribution of train and test PEs, we explore a possible mitigation, which flattens the distribution of PEs during training. We observe a consistent improvement of automatic metric scores for the APE-based vanilla encoder-decoder model, while the RoPE based decoder-only model remains mostly unaffected. In summary, our main contributions are: (a) a new approach to the detection and diagnosis of length issues in document-level MT, (b) a new variant of the SHAPE Kiyono et al. (2021) method, which improves the distribution of PEs during training, and (c) a confirmation that (perhaps for lack of an appropriate document-level evaluation tool) sentence-level MT remains a strong baseline in most settings.
2 Related work
2.1 Document-level MT
Previous attempts to incorporate more contextual information in MT models can be roughly categorized into two categories: context-augmented MT, also called Doc2Sent in (Sun et al., 2022), and document-level MT, also called Doc2Doc. Recent surveys of this field include (Popescu-Belis, 2019; Maruf et al., 2021; Castilho and Knowles, 2024).
Translation of discourse phenomena, such as lexical consistency, reference, and word sense disambiguation, requires inter-sentential context (Bawden et al., 2018; Wong et al., 2020; Fernandes et al., 2023). This has motivated the integration of extended (local) contexts in Doc2Sent models. Such approaches include concatenation-based methods (Tiedemann and Scherrer, 2017); architecture adaptations to process context in different components of the same encoder (Ma et al., 2020; Wu et al., 2023), in a dedicated encoder (Voita et al., 2018; Zhang et al., 2018), or via hierarchical attention networks (Miculicich et al., 2018; Maruf et al., 2019; Yin et al., 2021); cache-based methods using a short-term MT memory (Maruf and Haffari, 2018; Tu et al., 2018; Yang et al., 2019; Dobreva et al., 2020) and multi-pass decoding algorithms (Voita et al., 2019; Yu et al., 2020; Kang et al., 2020).
Translating sentence by sentence, even with augmented contexts, still fails to capture phenomena related to coherence and consistency (Fernandes et al., 2023), motivating Doc2Doc approaches to process documents as a whole. This can be done with concatenation-based methods (Tiedemann and Scherrer, 2017; Sun et al., 2022; Karpinska and Iyyer, 2023), along with sliding window attention (Zhuocheng et al., 2023; Liu et al., 2023) and group attention (Bao et al., 2021) to address the issue of quadratic complexity. Other strategies include focusing on improving training through data augmentation with a balanced length distribution (Sun et al., 2022) and richer context-dependent phenomena (Lupo et al., 2022a; Wu et al., 2024), or on better training strategies with multilingual denoising pretraining (Lee et al., 2022), adapted loss functions (Lupo et al., 2022b), and enriched positional encodings (Li et al., 2023; Lupo et al., 2023). Multiple methods have recently emerged for large language models (LLMs) (Wang et al., 2023), which also show a decline in translation quality as input length increases (Wang et al., 2024). [^2] They also confirmed the effectiveness of training LLMs on documents of varied sizes (similar to Sun et al. (2022)). These include a two-stage training recipe with the use of a monolingual corpus and high-quality parallel documents (Xu et al., 2024; Alves et al., 2024), and applying LLMs as post-editors (Koneru et al., 2024).
2.2 Extrapolating PEs
disabledisabletodo: disableIntroduce APE / RPE
Since self-attention is position-agnostic, PEs are used to provide position information in Transformer models. PEs embed the absolute token position (APEs) (Vaswani et al., 2017), or the relative distance between tokens (RPEs) (Shaw et al., 2018; Raffel et al., 2020; Press et al., 2022), with RoPE (Su et al., 2024) being the go-to approach in recent LLMs such as Llama2 (Touvron et al., 2023). Despite RPEs yielding better length extrapolation ability than APEs, both of them struggle to efficiently extrapolate input lengths beyond the predefined maximum training length (Dai et al., 2019; Chen et al., 2023; Peng et al., 2024; Zhao et al., 2024), motivating the development of input extension methods for PEs.
For APEs, SHAPE (Kiyono et al., 2021) offsets all indices in a sequence by some random values. Its authors show that this simple technique mimics the computation of RPEs at a much smaller cost and helps to improve the interpolation abilities of a vanilla encoder-decoder model, as measured by BLEU (Papineni et al., 2002) with long pseudo-documents. Our experiments confirm that this technique is effective using actual document contexts and a sounder experimental methodology, based on paired tests, and using COMET (Rei et al., 2020).disabledisabletodo: disableRead this one ! Sinha et al.’s (2022) experiments adopt a setting similar to ours, offsetting the absolute value of APEs’ input to evaluate their ability to capture relative distances between tokens. Their results, like ours, illustrate the lack of robustness of APEs and suggest that they overfit their training data.
For RPEs, especially RoPE, both position interpolation (PI) and position extrapolation methods have been proposed. PI methods interpolate positions to extrapolate context length directly during inference or through fine-tuning (Chen et al., 2023; Peng et al., 2024). The position extrapolation methods aim to extend context using documents that are shorter than the predefined maximum length. For example, RandPos (Ruoss et al., 2023) randomly maps position indices to a much larger interval with the original word order, and PoSE (Zhu et al., 2024) divides each training sequence into $N$ chunks and adjusts the position indices of every chunk except the first one by adding a uniformly sampled offset, within the scope of a predefined maximal length.
3 Methods and Metrics
3.1 Holistic Document-Level MT
disabledisabletodo: disabletake some part from taln2024-long-docdisabledisabletodo: disableintroduction for ”such questions” mentioned belowdisabledisabletodo: disableExplain why this is interesting - for instance WMT 24 - full context setting
Compared to sentence-based MT, holistic document-Level MT (Doc2Doc) possesses several appealing features, as it gives access to all the available textual context. This should enable the MT system to improve on global aspects pertaining for instance to coherence and cohesion.disabledisabletodo: disableAdd refs. However, Doc2Doc also introduces several new challenges compared to the Sent2Sent scenario:
  1. 1.
    in Doc2Doc, input texts are longer, causing a computational overhead due to the quadratic complexity of attention (Tay et al., 2022).
  2. 2.
    for longer inputs, attention weights are spread over a larger number of tokens (Herold and Ney, 2023); however, at each decoding step, most attention needs to remain concentrated on the corresponding local source context (Bao et al., 2021). This is in contrast with Doc2Sent, where sentence alignment is readily available.
  3. 3.
    decoding longer sequences increases the impact of search errors and of exposure bias (Ranzato et al., 2016). Beam search also becomes more difficult due to the input length.
  4. 4.
    output sentences may not always stand in one-to-one correspondence with source sentences, which complicates the computation of automatic metrics, which are designed to evaluate one-to-one mappings between hypotheses and references.
These differences motivate our main research questions, which we rephrase as: (a) For existing models, does Doc2Doc bring more benefits than disadvantages compared with Sent2Sent? (b) How do these results vary with the input document length? (c) Which methods and metrics can we use to automatically evaluate the impact of length differences?
3.2 Shades of BLEU
Answering such questions requires metrics for comparing holistic translations with sentence-based translations: as the number of segments produced by the former may differ from the number of source segments, a basic requirement is that they allow the evaluation of translation hypotheses with more (or fewer) sentences than the source (for quality estimation scores) and/or the reference (for reference-based metrics). However, most existing document-level MT approaches still rely on BLEU Papineni et al. (2002), despite its well-documented shortcomings (Callison-Burch et al., 2006; Reiter, 2018; Mathur et al., 2020; Dahan et al., 2024); or rather a variant dubbed d-BLEU by Liu et al. (2020). [^3] Hendy et al. (2023) also consider a variant of COMET Rei et al. (2022) while Zhuocheng et al. (2023) introduce d-ChrF, a document-level version of ChrF (Popović, 2015). We accordingly focus on BLEU in this section, noting that the same questions would need to be addressed with any metric relying on sentence-based surface comparison (e.g., METEOR Banerjee and Lavie (2005), TER Snover et al. (2006), BertSCORE (Zhang et al., 2020), PRISM (Thompson and Post, 2020), COMET (Rei et al., 2020), and many others). [^4] We choose to evaluate using the standard metrics, BLEU and COMET, rather than evaluation approaches specifically designed to test the use of increased context. This choice is motivated by the fact that the score differences we observe reveal a significant degradation in translation quality for longer documents, indicating greater problems than those targeted by finer-grained evaluation techniques.
BLEU is computed by counting, sentence by sentence, the number of $n$-grams (for $n\in[1\!:\!4]$) shared by each translation hypothesis and its human reference. These counts are aggregated and turned into frequencies, then averaged (geometrically) at the corpus level. Finally, a length penalty is applied to degrade the score when the cumulated length of the hypotheses is shorter than that of the references. BLEU is a corpus-level score that depends on sentence alignments. d-BLEU is also a global score but counts common $n$-grams at the document level. As a consequence, d-BLEU, which records matches for larger spans than BLEU, delivers higher scores, as the opportunities to match $n$-grams are greater for a wider window. [^5] This effect is well known, e.g. in (Koehn and Knowles, 2017, Figure 7), where BLEU increases when considering sentence groups of increasing lengths (at least for a certain length range), where we would expect a decrease, as the length is often linked to syntactic complexity and therefore to translation difficulty. We reproduce this observation in Figure 2. These two scores cannot be compared, and we contend that their shortcomings make them inappropriate for analysing length-related issues in MT.
An alternative to d-BLEU is to perform evaluation at the document level, rather than the corpus level. This can be implemented either as (a) calculating one BLEU score (with realignment) per document, then averaging at the corpus level or (b) calculating the equivalent of sentence-level BLEU scores (Lin and Och, 2004) but where each segment is a concatenated full document rather than a sentence. However, (a) counts matches at the sentence-level, which requires a realignment between translated and reference sentences and may introduce some measurement noise. Therefore, our experiments use method (b) to compute document-level scores, hereafter referred to as ds-BLEU scores.
Refer to captionRefer to captionRefer to caption
Figure 1: Top: probability of observing training position $i$ ($P(i)$) for a sentence of length $l=200$, with standard training ($k_{i}=0$, left) and with our uniform sampling scheme (right) for $M=512$. Middle: original, unifPE, and SHAPE $P(i)$ for training set TED-G and $M=2048$. Bottom: original, unifPE and SHAPE $P(i)$ for TED-U and $M=2048$.
3.3 Evaluating Length Issues in MT
Another recurring methodological caveat with length-related evaluation is related to the way scores are compared. For instance, in (Sun et al., 2022, Figure 1) BLEU scores are reported for buckets of sentences of varying lengths in a plot which suggests that performance increases with length (up to a certain extent). Such vizualisations are misleading, as global BLEU scores should only be compared when measured with the same corpus.
What we propose instead is to compare matching automatic translation scores for a set of inputs $\mathcal{S}=\{s_{1}\dots s_{T}\}$, systematically varying the translation models $M$ in $\{M_{1}\dots M_{N}\}$ and the length of the translation window $W\in\{W_{1}\dots W_{K}\}$. For each pair of settings, we can perform a paired t-test for the average score difference and decide whether two configurations $(M_{i},W_{k})$ and $(M_{j},W_{l})$, each associating a system and a length, are statistically different, and if so, which of the two is the best.
In our experiments, we consider two ways of presenting $\mathcal{S}$: (a) at the document level, where each $s_{i}$ is a document and the evaluation is the ds-BLEU score introduced in Section 3.2, [^6] We use SacreBLEU (Post, 2018) with signature:
nrefs:1||||case:mixed||||eff:no||||tok:13a||||smooth:exp||||version:2.4.0; the parameter eff is set to yes for ds-BLEU.
and (b) at the sentence level, where each $s_{i}$ is a sentence and the associated metric is COMET (Rei et al., 2020). [^7] Using the library https://github.com/Unbabel/COMET with the default model wmt22-comet-da. For (b) we need to realign translation hypotheses with their references. This can be performed with the method of Wicks and Post (2022), [^8] Junczys-Dowmunt (2019)’s approach includes a set of tags that constrain input and output to have the same number of sentences, see also (Li et al., 2023). or with that of Matusov et al. (2005), [^9] https://www-i6.informatik.rwth-aachen.de/web/Software/mwerSegmenter.tar.gz/. which has long been used for evaluating speech translation systems, and which we adopt. [^10] The per-sentence COMET scores are averaged at the document level to be associated with document lengths, or at the corpus level to assess global translation quality. Variations in configurations $(M,W)$ are obtained by changing the translation engine and the length of input source texts. In all cases, score comparisons are performed on identical source texts.
disabledisabletodo: disableAlso position: check
The same technique is also used to measure the impact of the position within a document on translation quality. The question we study is whether quality remains constant across a document, or whether it tends to decrease when sentences are processed at higher position indices. For this, we consider groups of sentences translated at varying starting positions with multiple systems and compare the differences between COMET scores with a paired difference test. Details regarding the corpus and window sizes are given in Section 5.1.
4 Manipulating the Distribution of PEs
A basic requirement for document-level systems is that they should be trained, or at least fine-tuned, with long text inputs, ideally with complete documents. Using the empirical document length distribution may however not be ideal, as it yields very skewed distributions of PEs where small position indices are over-represented. We discuss two approaches to obtain more balanced distributions.
4.1 Distribution of PEs
disabledisabletodo: disableExplain there are 2 ways: changing the pseudo length; shifting PEs
A training sequence of length $l$ yields examples for all indices in $\{1,\dots,l\}$. For a complete corpus, position index $1$ will be observed for all inputs, while the last index of the longest sequence will likely only be observed once. Training with the “natural” distribution of document lengths is therefore likely to overfit to smaller position indices while underfitting to larger ones, hindering the ability to handle long texts or extrapolate to lengths unseen in training (Peng et al., 2024; Zhu et al., 2024).
A first way to improve the distribution of token positions seen in training is to increase the representation of long documents in the training data while keeping a good balance with shorter ones (Bao et al., 2021; Sun et al., 2022). This is easy to do in our controlled setting (see Section 5.1). As our experiments show, this significantly improves automatic scores for the context lengths seen during training. An alternative, which allows us to better study the effect of PE distributions in training, is to directly manipulate the indices (for a fixed length distribution). The unifPE algorithm, introduced below, is one way to achieve this.
4.2 Uniform SHAPEs (unifPE)
We assume a training set of texts $s_{1}\dots s_{N}$ of respective lengths $l_{1}\dots l_{N}$, and a maximum model length of $M$, with $\forall i,M>l_{i}$. Training with text $s_{i}$ creates training samples for positions $i$ in $[1\!:\!l_{i}]$. For the whole corpus, positions from $1$ (observed $N$ times) to $l_{max}=\mathrm{max}_{i=1\dots N}(l_{i})$ are observed, with larger indexes being less trained than smaller ones. Positions indices in $[l_{max}\!:\!M]$ are never observed. We wish to make the training PE distribution more even, so that all positions in $[1\!:\!M]$ are equally well-trained, which should also help to extrapolate PEs for indices larger than $l_{max}$.
This can be achieved by shifting the starting index of every $s_{i}$ by some offset $k_{i}$, making it possible to train with PEs in $[1+k_{i}\!:\!l_{i}+k_{i}]$. How should $k_{i}$ be chosen? Randomly choosing $k_{i}=0$ or $k_{i}=l_{i}$ with probability $1/2$ makes the probability of observing any index in $[1\!:\!2l_{i}]$ equal to $1/2$. This can be generalised to choose $k_{i}$ with uniform probability $1/m$ among $\{0,l_{i},\dots,(m-1)*l_{i}\}$, with $m=\lfloor M/l_{i}\rfloor$. However, doing so implies that indices in $[m*l_{i}:M]$ are never observed. We compensate for this as follows: before sampling $k_{i}$, we modify the set of possible shifts by adding $r_{i}=M-m*l_{i}$ to all values larger than a random index $l\in[1\!:\!m]$. In other words, $k_{i}$ is sampled from $\{j*l_{i}+r^{\prime}_{i,j},j=0\dots m-1\}$, with $r^{\prime}_{ij}=0$ if $j<l$ and $r_{i}$ otherwise. Sampling $k_{i}$ independently for each text $s_{i}$ in each training batch ensures that all indices are uniformly represented. A formal description of unifPE is given in Algorithm 1. Figure 1 illustrates the difference between always starting at position $1$ ($\forall i,k_{i}=0$) and using our unifPE strategy.
This approach is reminiscent of SHAPE (Kiyono et al., 2021); while SHAPE chooses the offset $k_{i}$ uniformly at random in a fixed interval to simulate relative PEs, which reduces the frequency of small position indices, we sample $k_{i}$ non-uniformly to ensure that all indexes are equally represented in training.
5 Experimental Settings
5.1 Datasets
For our experiments, we prepare multiple sets of parallel pseudo-documents based on the EN–FR part of the TEDtalks corpus (Cettolo et al., 2012).
{floatrow}
TED-full TED-G TED-U
train dev train dev train dev
Count 1831 19 15625 160 10582 106
Length 2915 2861 341 339 504 512
sent 256 512 768 1024 1200 1600 2048 doc
Count 5103 503 261 184 142 123 100 80 52
Length 23 233 450 638 827 955 1175 1468 2259
Table 1: Left: Statistics of the TED talks training and dev sets. Right: Statistics of the TED talks test sets from IWSLT tst2014, tst2015, tst2016 and tst2017. ‘Count’ denotes the number of parallel pseudo-documents, ‘Length’ denotes the average length of source (i.e. English) pseudo-documents (in NLLB tokens).
$l_{max}$ 2014 2015 2016 2017
NLLB sent 45.1 (0.97) 43.9 (0.98) 41.7 (1.00) 41.8 (1.00)
256 33.9 (0.82) 35.4 (0.84) 33.3 (0.86) 33.5 (0.87)
512 14.6 (0.44) 16.0 (0.56) 15.2 (0.52) 13.8 (0.49)
768 7.3 (0.27) 7.9 (0.32) 10.0 (0.46) 6.7 (0.27)
1024 8.8 (0.56) 7.4 (0.51) 7.5 (0.50) 6.5 (0.48)
TowerBase sent 43.4 (0.98) 42.9 (0.99) 39.7 (1.00) 38.7 (1.00)
256 44.0 (0.96) 42.8 (0.98) 40.9 (1.00) 39.4 (1.00)
512 42.9 (0.96) 39.8 (0.98) 39.9 (1.00) 40.6 (1.00)
768 39.6 (0.98) 39.0 (0.97) 38.1 (0.99) 39.9 (1.00)
1024 38.5 (0.98) 33.1 (0.99) 35.4 (1.00) 35.4 (0.98)
1200 37.4 (0.92) 35.5 (0.98) 36.2 (1.00) 35.6 (0.98)
1600 33.3 (0.96) 34.9 (0.96) 26.7 (0.94) 31.0 (0.97)
2048 24.0 (0.97) 27.7 (0.95) 27.2 (0.96) 23.5 (0.87)
Table 2: ds-BLEU scores (and brevity penalty) for NLLB200-distilled-600M and TowerBase-7b.
NLLB FT-U Unif-U SHAPE-U FT-G Unif-G SHAPE-G
sent-256 9.2 0.8 -2.1 -2.5 0.4 -0.7 -1.7
256-512 19.1 - -0.4 1.4 - - 2.1
512-768 6.9 - -0.6 - 5.9 2.5 5.3
768-1024 - 0.5 - - 7.2 3.0 3.7
1024-1200 2.2 3.5 1.9 4.1 4.0 3.3 4.1
1200-1600 - 6.8 6.5 5.4 5.8 5.5 4.9
1600-2048 1.9 5.2 4.5 3.1 4.2 5.7 6.0
sent-256 16.7 3.5 1.7 1.3 2.7 2.1 1.9
256-512 20.7 - -0.4 2.4 0.6 - 4.2
512-768 5.6 - - - 11.6 7.2 8.1
768-1024 5.2 2.3 0.8 - 10.4 7.6 7.8
1024-1200 - 7.4 4.3 6.9 3.8 4.7 4.6
1200-1600 6.1 9.6 13.4 8.3 5.4 5.5 5.6
1600-2048 - 5.1 5.0 5.9 3.9 5.6 5.0
Tower FT-U Unif-U SHAPE-U FT-G Unif-G SHAPE-G
sent-256 - - - - - - -
256-512 - 0.9 0.8 0.8 0.6 0.6 0.5
512-768 - - - - 0.6 - -
768-1024 3.4 - 1.0 1.2 1.7 1.2 2.1
1024-1200 - - - - - - -
1200-1600 4.7 1.7 2.1 - 1.6 2.0 1.5
1600-2048 5.9 7.5 6.5 7.3 8.1 7.8 7.3
sent-256 3.9 2.3 2.4 2.3 2.3 2.3 2.2
256-512 - 0.4 - - 0.2 0.3 0.3
512-768 - - - 0.5 0.3 - -
768-1024 2.9 1.0 0.5 - 1.1 0.8 1.2
1024-1200 - - 1.0 0.9 - - -
1200-1600 6.2 1.7 1.8 - - 1.9 1.8
1600-2048 8.7 10.0 8.9 9.1 11.0 10.2 9.2
Table 3: Average differences evaluated on ds-BLEU (top) of full TED talks and on 100×100$\times$100 ×COMET (bottom) of realigned parallel sentences, between translations in increasing context size, for NLLB (left) and TowerBase (right) models. U and G respectively denote TED-U and TED-G. A positive value means that shorter segments result in higher scores than longer ones. Text in olive for p-values $>0.01$. - for p-values $>0.05$.
Training and validation sets
Our training set consists of pseudo-documents from both the training and validation splits of IWSLT-2016.111111https://wit3.fbk.eu/2016-01 Our goal is to simulate real corpora of parallel documents with source documents shorter than a certain length $l_{max}$ – using $l_{max}=1024$. We split all document pairs whose source side is longer than 1024 tokens into fragments. [^12] All statistics counted in tokens use the tokeniser of NLLB (Costa-jussà et al., 2024). For each document pair, we iterate the following procedure: (1) sample a maximum pseudo-document length $l^{\prime}_{i}$ following the same Gaussian-like length distribution as the full TED talks with $l^{\prime}_{i}<l_{max}$, (2) concatenate consecutive sentence pairs up to $l^{\prime}_{i}$ to form a training pseudo-document $s_{i}$. The resulting distribution of document lengths is displayed in Figure 3 in Appendix A.3. The development set is built similarly, using document pairs from IWSLT tst2010 and tst2011. We denote these training datasets as TED-G (G for Gaussian). As discussed in Section 4, we consider another dataset generation strategy, which produces a more balanced length distribution, for which we do as above but we sample uniformly: $l^{\prime}_{i}\sim U(128,l_{max})$. [^13] Short pseudo-documents continue to be slightly over-represented, because the last pseudo-document in any given talk is often strictly shorter than the desired length $l^{\prime}_{i}$. Fine-tuning with the resulting TED-U corpus allows us to contrast two distributions with differences in document length.
Test sets
To evaluate MT systems for their ability to handle documents of varying sizes and extrapolate beyond the training samples, we build a series of test sets of increasing document lengths. For each document in IWSLT tst2014, tst2015, tst2016 and tst2017, we accumulate consecutive sentence pairs into parallel pseudo-documents such that all resulting source texts have a length close to $l_{max}$, with $l_{max}\in\{256,512,1024,1200,1600,2048\}$. [^14] At the end of each talk, we concatenate the last parallel sentences into the last pseudo-document if they are shorter than $50$ to avoid exceedingly short parallel sequences. Contrarily to training sets, test sets are homogeneous in length. Statistics are in Table 1 with more details in Appendix A.3. Evaluation is always performed with complete original talks, after concatenating and aligning all the corresponding parts.
5.2 Models
We used the unifPE algorithm to fine-tune two pre-trained MT systems that were not trained with TED talks. As unifPE is designed for APEs, we considered NLLB200-distilled-600M [^15] https://huggingface.co/facebook/nllb-200-distilled-600M or NLLB for short (Costa-jussà et al., 2024) as a representative encoder-decoder model based on APEs. NLLB is a 12-layer encoder-decoder multilingual MT model pre-trained on 200 languages. We used the HuggingFace implementation, which relies on sinusoidal APEs (Vaswani et al., 2017). We also perform fine-tuning with SHAPE for comparison. We refer to the specific MT systems with respect to their fine-tuning method (FT, unifPE or SHAPE), backbone model (e.g. NLLB) and training corpus (U for TED-U or G for TED-G. More precisely, we denote MT systems trained on TED-U (resp. TED-G) as FT-NLLB-U (resp. FT-NLLB-G), Unif-NLLB-U (resp. Unif-NLLB-G) when fine-tuning with unifPE, and SHAPE-NLLB-U (resp. SHAPE-NLLB-G).
We also experiment with an LLM-based architecture, TowerBase-7B [^16] https://huggingface.co/Unbabel/TowerBase-7B-v0.1 (Alves et al., 2024) (TowerBase for short), derived from Llama2 (Touvron et al., 2023) using translation-related tasks. TowerBase uses RoPE (Su et al., 2024) to encode RPEs. As mentioned by Peng et al. (2024), they nonetheless encode some form of APE signal in some dimensions, and may therefore be also mildly impacted by the PE training distribution. We refer to the models based on TowerBase as FT-Tower-U (resp. FT-Tower-G), Unif-Tower-U (resp. Unif-Tower-G) and SHAPE-Tower-U (resp. SHAPE-Tower-G) the model fine-tuned on TED-U (resp. TED-G) with original PEs, unifPE or SHAPE.
Both backbone models were pretrained with large amounts of EN–FR data; we focus exclusively on the EN into FR direction. disabledisabletodo: disableCheck directionDetails on fine-tuning and decoding parameters can be found in Appendix A.4.
6 Results and Analyses
disabledisabletodo: disableput all results about ratio7 in an archiv
6.1 Length Issues
We report the ds-BLEU (Table 2) and COMET (Appendix, Table 7) scores of the pretrained models NLLB and TowerBase for multiple test sets, varying the average input segment lengths from one sentence to the maximum input length used in training. [^17] As explained in Section 3.2, these COMET scores require the realignment of target sentences with the reference.disabledisabletodo: disableCOMET scores in appendix ? For NLLB, we observe a drop of around $10$ ds-BLEU points and about $0.2$ COMET points when translating test sets of $l_{max}=256$ instead of isolated sentences. Scores and their associated brevity penalties (BPs) only get worse with larger context lengths. For TowerBase, the decrease in BLEU is more progressive, with a sharp decline for all test sets for $l_{max}>1024$. The related COMET scores plummet immediately with a context size of $256$. Even though TowerBase is based on Llama2, which accepts inputs up to $4096$ tokens, the continued pretraining that was used mostly uses isolated sentences, which introduces an inductive bias affecting its ability to translate long texts.
As expected, document-level fine-tuning (DLFT) has a strong positive impact (see Appendix, Table 11). However, the length issues remain.
TED-U TED-G FT Unif SHAPE
 FT vs Unif  FT vs SHAPE  FT vs Unif FT vs SHAPE U vs G U vs G U vs G
sent 3.3 (0.00) 4.0 (0.00) 1.2 (0.00) 2.4 (0.00) - -2.1 (0.00) -1.6 (0.00)
256 - 0.7 (0.00) - - - -0.6 (0.01) -0.7 (0.00)
512 -0.5 (0.01) 1.7 (0.01) - 2.3 (0.00) -0.7 (0.00) -0.4 (0.01) -
768 -0.8 (0.00) 2.7 (0.00) -3.7 (0.00) - 5.5 (0.00) 2.6 (0.00) 4.5 (0.00)
1024 - 1.6 (0.00) -7.8 (0.00) -1.8 (0.04) 12.2 (0.00) 5.1 (0.00) 8.8 (0.00)
1200 -2.3 (0.00) 2.3 (0.02) -8.5 (0.00) - 12.7 (0.00) 6.5 (0.00) 8.8 (0.00)
1600 -2.6 (0.01) - -8.8 (0.00) -2.6 (0.01) 11.8 (0.00) 5.6 (0.00) 8.3 (0.00)
2048 -3.3 (0.00) - -7.3 (0.00) - 10.7 (0.00) 6.8 (0.00) 11.3 (0.00)
sent 1.9 (0.00) 2.9 (0.00) 0.9 (0.00) 1.4 (0.00) - -0.9 (0.00) -1.4 (0.00)
256 - 0.8 (0.00) 0.4 (0.00) 0.7 (0.00) -0.8 (0.00) -0.5 (0.01) -0.9 (0.00)
512 -0.6 (0.02) 2.9 (0.00) - 4.3 (0.00) -0.4 (0.03) - -
768 -0.7 (0.00) 4.7 (0.00) -4.7 (0.00) - 11.2 (0.00) 7.2 (0.00) 7.3 (0.00)
1024 -2.2 (0.00) 3.3 (0.00) -7.5 (0.00) -1.8 (0.02) 19.3 (0.00) 14.0 (0.00) 14.2 (0.00)
1200 -5.3 (0.00) 2.7 (0.02) -6.5 (0.00) - 15.7 (0.00) 14.5 (0.00) 11.9 (0.00)
1600 - - -6.4 (0.00) - 11.5 (0.00) 6.6 (0.00) 9.2 (0.00)
2048 - 2.1 (0.04) -4.7 (0.00) - 10.4 (0.00) 7.2 (0.00) 8.3 (0.00)
Table 4: Average difference (and p-values) in ds-BLEU (top) evaluated on full TED talks and 100×100$\times$100 ×COMET (bottom) evaluated on realigned sentences for NLLB. Left and middle: paired comparison between fine-tuning with the original PEs (FT), unifPE (Unif) and SHAPE on TED-U and TED-G respectively. Right: differences between fine-tuning on TED-U (U) and TED-G (G). - for p-values $>0.05$. Bold values when the two metrics disagree on significativity.
Length Bias
We performed paired comparisons for the translation of our test sets with increasing text lengths for each MT system as presented in Section 3.3. Results are given in Table 3, where a positive difference (e.g. $9.2$ for Nllb in line “sent-256”) means that the translation of shorter segments (here: sentences) yield better scores than that of longer ones ($256$ tokens). Scores in the same column are comparable. Except for a handful of configurations, translating longer texts is never better than translating short ones. We conclude that in our experimental settings, the disadvantages associated with long inputs (Section 3.1) overwhelm the benefits of a complete context. These length issues result in large score degradations and are not easily fixed by simple manipulation of PEs. We also observe that results obtained with COMET and ds-BLEU sometimes disagree. These cases are rare, though, suggesting that our results are robust.
Document-level Tuning with unifPE
Again using the paired comparison methodology, we compare the performance of DLFT with original PEs, unifPE and SHAPE. As shown in the left and middle parts of Table 4, fine-tuning using unifPE leads to steady improvements in translation scores for all test lengths, especially for systems fine-tuned with the unbalanced corpus (TED-G). The only exception is for sentence-level translations, which remain marginally better using standard DLFT than with unifPE. In contrast, SHAPE only improves DLFT performance on the TED-G corpus and for translation windows greater than $1024$ tokens, due to the under-representation of small position indices during training, as shown in Figure 1. As Tables 11, 12, 13, 14 and 15 show, these improvements remain moderate, and the length issues continue to strongly impact translation scores, especially for test documents of $1024$ tokens or more.
NLLB FT-U Unif-U SHAPE-U FT-G Unif-G SHAPE-G
$p_{0}$-$p_{1}$ 12.7 - - 4.0 1.6 3.0 5.0
$p_{1}$-$p_{2}$ 7.2 - - -2.0 1.9 2.4 -
$p_{2}$-$p_{3}$ 2.9 1.0 -1.3 - 2.4 - -2.1
$p_{3}$-$p_{4}$ 7.2 5.5 4.6 7.3 26.1 10.7 13.9
$p_{4}$-$p_{5}$ 3.9 - - 8.3 4.7 4.3
$p_{5}$-$p_{6}$ 3.3 31.6 27.1 19.5 6.1 15.4 15.3
Tower FT-U Unif-U SHAPE-U FT-G Unif-G SHAPE-G
$p_{0}$-$p_{1}$ 1.2 - - - - - -
$p_{1}$-$p_{2}$ 0.6 0.6 -
$p_{2}$-$p_{3}$
$p_{3}$-$p_{4}$ 4.9 1.1 0.6 1.4 1.1 1.3 1.1
$p_{4}$-$p_{5}$ 1.7 1.4 2.1 1.7 2.0 1.8 2.2
$p_{5}$-$p_{6}$ 26.3 24.5 26.0 25.2 27.1 26.4 27.0
Table 5: Average difference of 100×100$\times$100 ×COMET-score evaluated on $794$ sentence pairs, translated at different positions (e.g. $p_{0}$ and $p_{1}$ with $p_{0}<p_{1}$) by NLLB-based systems (top) and TowerBase-based systems (bottom). Olive text for p-values $>0.01$. - for p-values $>0.05$.
For TowerBase, unifPE does not yield any significant difference with standard DLFT, and SHAPE occasionally delivers slight improvements (see Appendix, Table 16), likely because this model relies on RPEs. From these comparisons, we conclude that unifPE partly resolves length issues for nllb, but hardly changes the situation for TowerBase.
Impact of Data Distribution
In the right part of Table 4, we evaluate the impact of the length distribution during fine-tuning for NLLB: the balanced distribution (TED-U) slightly but consistently underperforms the use of TED-G for short documents (fewer than $512$ tokens), a trend that is reversed for longer documents with strong improvement (over $768$ tokens). Manipulating the distribution of PEs with unifPE reduces the gap between the two fine-tuning corpora and makes the model more robust to document lengths rarely observed (or even unobserved) during fine-tuning. This analysis again reveals small differences between using ds-BLEU and COMET scores: in nine cases out of $56$ comparisons (marked in bold), one metric detects a difference that is non-significant for the other.
6.2 Position Bias
To investigate potential translation issues related to large position indices, we collected the $794$ sentences that come from the final part of long talks and for which varying the window length also varied the position index. For each of them, we have seven translations, corresponding to positions $\{p_{0}^{j},\dots,p_{6}^{j}\},j\in\{1,\dots,794\}$. The average values for $\{p_{0}^{j},\dots,p_{6}^{j}\}$ are $\{p_{0},\dots,p_{6}\}=[66,173,262,335,585,779,1477]$. For this subset of sentences, we performed a paired t-test to compare the impact of the position on the translation score (using COMET as the only metric). We observe in Table 5 that in almost all comparisons but three, a small position index is preferable to a larger one. This suggests that one of the main challenges faced by Doc2Doc with large context lengths is to control the quality degradation for the final parts of the input text. Here again, unifPE slightly mitigates the problem for NLLB models compared with original PEs and SHAPE, but no such improvement is observed for TowerBase.disabledisabletodo: disableWhy is tower different?
256 512 768 1024 1200 1600 2048
NLLB 0.04 0.35 0.49 0.66 0.64 0.74 0.81
FT-U 0.01 0.03 0.08 0.09 0.13 0.26 0.44
Unif-U 0.01 0.03 0.07 0.10 0.20 0.34 0.32
SHAPE-U 0.03 0.08 0.11 0.20 0.36 0.39 0.46
FT-G 0.01 0.03 0.11 0.31 0.40 0.57 0.69
Unif-G 0.01 0.04 0.16 0.20 0.27 0.25 0.36
SHAPE-G 0.02 0.05 0.12 0.17 0.21 0.24 0.28
256 512 768 1024 1200 1600 2048
Tower 0.01 0.05 0.13 0.30 0.29 0.45 0.64
FT-U 0.01 0.05 0.10 0.15 0.13 0.23 0.59
Unif-U 0.02 0.05 0.09 0.15 0.16 0.28 0.61
SHAPE-U 0.02 0.05 0.08 0.15 0.17 0.21 0.59
FT-G 0.02 0.05 0.10 0.15 0.19 0.26 0.62
Unif-G 0.02 0.07 0.10 0.16 0.20 0.28 0.64
SHAPE-G 0.02 0.05 0.07 0.17 0.17 0.27 0.62
Table 6: Percentage of pseudo-documents among IWSLT tst2014-2017 in which $10$-gram repetition is detected in the translation given by NLLB-based (top) and TowerBase-based models (bottom).
6.3 Repeated $n$-grams in Translation
One obvious problem with holistic translations produced by Nllb is the generation of outputs that are too short. A closer look at translation outputs also reveals that outputs contain many instances of repeated texts, usually occurring in the final part of the translation. To quantify this problem, we compute the percentage of translations of pseudo-documents in which the repetition of a long $n$-gram (with $n\geq 10$) is detected. Detailed results are given in Table 6. For all systems and fine-tuning strategies, the percentage of repetitions increases with the length, a problem that seems (for large text lengths) slightly more severe for TowerBase, which has a much better BP, than Nllb.
7 Conclusion
disable,color=cyan!30disable,color=cyan!30todo: disable,color=cyan!30More criticisms of BLEU
In this work, we have studied the ability of current MT architectures to handle long input texts, ideally entire documents, and to translate them holistically. Our analyses are based on systematic comparisons of translation outputs computed with varying input lengths, which are then evaluated with two automatic metrics. They consistently show that, even when the test document lengths match that of the training set and remain within the model limits, the translation scores tend to decrease with the source length, a degradation that mostly impacts sentences occurring far from the beginning of the document. We also show that manipulating the training distribution of lengths or PEs has a positive effect for APE-based models, which vanishes in RoPE-based models like TowerBase.disabledisabletodo: disableMake sure RPE is defined These results finally confirm the robustness of sentence-level baselines. They hint at the need to improve existing models to truly benefit from the potential of document-level MT, for instance by constraining the attention mechanism to simulate a form of sentence alignment, by improving the memorization capacities of existing architectures, or by ensuring that the generation algorithm does not eventually get trapped in repetition loops. These are some of the directions we wish to explore in future work.disabledisabletodo: disablePerhaps: more citations for this disable,color=cyan!30disable,color=cyan!30todo: disable,color=cyan!30Future: the errors beyond repetitions and missing material disabledisabletodo: disablethere should not be TEDtalks in NLLB training set, as the NLLB papers used IWSLT TED talks as benchmark to show the model performance, where is the only mention os TED talks.
8 Limitations
The empirical observations reported in this paper are based on one single language direction,disabledisabletodo: disablecheck that this is clear and one domain (TEDtalks). This experimental design is motivated by (a) the fact that French-English is considered an easy pair for MT, with large sets of parallel training data; (b) TEDtalks data are a standard benchmark for document-level MT, and crucially contain very long parallel documents, allowing us to implement our evaluation methodology on a large range of length values. Furthermore, these datasets are not included in the training data of our models. We contend that the length issues observed in these favorable conditions for two representative systems would only be worse for more difficult or less-resourced language pairs.
9 Ethics Statement
This study has been performed with standard benchmarks and open-weight models. We do not see any ethical problems with this work.
10 Carbon Impact Statement
The experiments were conducted on a private infrastructure using a single A100 SXM4 GPU, with a carbon efficiency of $0.432$ kgCO2eq/kWh. The average time required for fine-tuning and checkpoint selection was $14.21$ hours for the six NLLB models, and $5.6$ hours for the Towerbase models. The average emissions are estimated to be $2.45$ kgCO2eq for NLLB-based models and $0.97$ kgCO2eq for models derived from TowerBase, with no offset applied. These estimations were based on the Machine Learning Impact calculator [^18] https://mlco2.github.io/impact#compute (Lacoste et al., 2019).
Acknowledgments
This work was supported by the French national agency ANR as part of the MaTOS project. [^19] http://anr-matos.fr/ Rachel Bawden was also partly funded by her chair position in the PRAIRIE institute funded by ANR as part of the “Investissements d’avenir” programme under reference ANR19-P3IA-0001. The authors are grateful to the anonymous reviewers for their insightful comments and suggestions and to Paul Lerner for his review and feedback on a preliminary draft of this work.
Appendix A Appendix
A.1 The unifPE Algorithm
The unifPE algorithm briefly described in Section 4.2 is formalised in Algorithm 1.
Data: $l_{i}$: The input length
Data: $M$: The target max context length
Data: $\operatorname{List_{p_{k}}}$: the distribution of $p_{k}$ for each offset $k$ in $[0,M-l_{i}]$
$\operatorname{List_{p_{k}}}\longleftarrow\text{Initialized to $0$ for each element}$
$m\longleftarrow\lfloor M/l_{i}\rfloor\text{ nb.\ of possible non-zero }p_{k}$
$R_{n}\longleftarrow\text{the remainder of $M$ divided by }l_{i}$
$p_{0}\longleftarrow\frac{1}{m}\text{the probability of each non-zero $p_{k}$ }$
if $M<2l_{i}$ then
       $\operatorname{List_{p_{k}}}\longleftarrow p(k^{\prime}=0)=1\text{ i.e.\ }p_{k=0}=1$
else
      $k^{*}\longleftarrow\text{a random integer in $[0,m)$}$
      for $k\in[0,M-l_{i}]$ do
             if $k\%l_{i}==0$ and $k<k^{*}$ then
                   $\operatorname{List_{p_{k}}}\longleftarrow p(k^{\prime}=k)=p_{0}$
             end if
            if $(k-R_{n})\%l_{i}==0$ and $k^{*}<k\leq M-l_{i}$ then
                   $\operatorname{List_{p_{k}}}\longleftarrow p(k^{\prime}=k)=p_{0}$
             end if
       end for
end if
return $\operatorname{List_{p_{k}}}$
Algorithm 1 unifPE: the pseudo-uniform position indices mapping algorithm.
A.2 A Call for Correctly Using BLEU Scores
As illustrated in Figure 2, d-BLEU and ds-BLEU are always larger than BLEU. When BLEU decreases due to the degradation of translation quality, d-BLEU remains stable because of the higher probability to find $n$-gram matches in longer sequences. In contrast, ds-BLEU consistently decreases when BLEU diminishes, as it applies a macro-average to compute the corpus-level score, which is more sensitive to the translation quality of each document than d-BLEU. Therefore, d-BLEU, ds-BLEU and BLEU are not comparable and d-BLEU is not suitable for analysing length issues in document-level evaluation of MT.
Refer to captionRefer to caption
Figure 2: BLEU, ds-BLEU and d-BLEU scores for IWSLT tst2015, translating and evaluating pseudo-documents of increasing lengths $[256,512,768,1204]$, using FT-NLLB-U (top) and Unif-Tower-U (bottom). Note that d-BLEU is computed for pseudo-documents while ds-BLEU is computed for concatenated full talks.
Refer to captionRefer to caption
Figure 3: Source document length distribution in the training set of TED-G (top) and TED-U (bottom).
A.3 Data Statistics and Other Details
Full data statistics are given in Tables 8 and 9. All the full TED talks in our corpora start with the title, then the description and the talk before being split into pseudo-documents. <description> and <title> tags are removed. When preparing our training and validation sets TED-U and TED-G (see Section 5.1), if concatenating the last sentence pair $(x_{n},y_{n})$ into the current pseudo-document pair exceeds $l_{max}$, $(x_{n},y_{n})$ will yield a single parallel sequence, to respect the maximum length $l_{max}$. The length distribution is illustrated in Figure 3.
2014 2015 2016 2017
NLLB sent 84 85 85 84
256 68 69 68 66
512 49 47 47 46
768 43 42 40 41
1024 36 37 36 36
TowerBase sent 84 85 85 85
256 80 81 82 80
512 79 80 82 80
768 78 78 80 80
1024 76 73 78 76
1200 73 74 77 76
1600 70 72 65 68
2048 52 63 65 57
Table 7: 100×100$\times$100 ×COMET scores for NLLB (top) and TowerBase (bottom).
count mean min max
TED-full train 1831 2915 /3515 56 /62 8706 /9706
dev 19 2861 /3460 680 /867 6076 /7590
TED-G train 15625 341 /411 3 /2 1022 /1460
dev 160 339 /410 12 /17 959 /1203
TED-U train 10582 504 /608 3 /1 1020 /1527
dev 106 512 /620 42 /41 991 /1276
Table 8: Statistics of the TED talks training and dev sets. ‘count’ denotes the number of parallel pseudo-documents. ‘mean’, ‘min’ and ‘max’ represent the average, minimum and maximum lengths of English/French pseudo-documents respectively, in NLLB tokens.
2014 2015 2016 2017
$l_{max}$ count min max mean count min max mean count min max mean count min max mean
EN sent 1335 2 112 23 1104 2 119 23 1185 1 151 24 1479 2 162 23
256 129 65 286 234 107 71 325 234 123 61 255 232 144 65 271 234
512 68 85 511 443 56 53 510 447 63 56 511 454 74 73 511 456
768 48 116 767 628 40 86 766 626 45 104 767 635 51 57 767 662
1024 37 83 1022 815 30 68 1023 835 35 115 1023 817 40 65 1023 844
1200 32 54 1218 942 26 71 1198 963 31 73 1216 922 34 125 1203 992
1600 26 135 1597 1160 24 114 1599 1043 23 191 1616 1243 27 229 1635 1250
2048 20 569 2091 1507 16 176 2072 1565 21 247 2046 1361 23 65 2045 1467
doc 15 995 4116 2010 12 1256 3359 2086 13 842 3366 2199 12 1909 3722 2812
FR sent 1335 2 158 28 1104 2 145 27 1185 1 180 29 1479 2 211 27
256 129 80 380 295 107 85 355 282 123 69 345 276 144 78 375 275
512 68 106 717 559 56 62 679 540 63 70 672 539 74 83 737 535
768 48 142 1065 792 40 102 1009 755 45 112 985 755 51 68 1083 776
1024 37 100 1436 1027 30 64 1349 1007 35 125 1314 970 40 80 1431 990
1200 32 61 1641 1188 26 85 1577 1162 31 93 1511 1096 34 134 1714 1164
1600 26 173 2188 1462 24 156 2116 1259 23 209 2074 1477 27 279 2261 1466
2048 20 657 2613 1901 16 218 2602 1889 21 280 2626 1617 23 80 2714 1721
doc 15 1289 4983 2534 12 1609 4013 2518 13 1004 4179 2613 12 2473 4464 3299
Table 9: Statistics of the test sets based on talks from IWSLT tst2014, tst2015, tst2016 and tst2017 (see Section 5.1). ‘count’ refers to the number of parallel pseudo-documents. ‘mean’, ‘min’ and ‘max’ denote the average, minimum and maximum lengths of the source (i.e. English, top) or the reference (i.e. French, bottom) pseudo-documents. All lengths are counted in NLLB tokens.
NLLB FT-U Unif-U FT-G Unif-G
sent 0 0 0 0 0
256 557 6 6 6 6
512 1231 5 9 12 11
768 1618 6 10 250 280
1024 886 53 34 491 486
1200 1179 437 207 576 675
1600 352 465 657 789 840
2048 456 644 843 801 1089
Tower FT-U Unif-U FT-G Unif-G
sent 0 0 0 0 0
256 79 3 8 2 3
512 58 2 2 3 2
768 65 4 3 3 5
1024 45 17 21 19 22
1200 107 19 17 13 19
1600 91 73 50 40 54
2048 151 94 84 66 98
Table 10: Number of empty alignments across all the $5,103$ sentences in our test sets for NLLB (top) and TowerBase (bottom) models.
A.4 Experimental Settings
This section presents detailed experiment settings for fine-tuning NLLB and TowerBase.
For NLLB, we fine-tuned the pretrained model with learning rate $5\mathrm{e}{-4}$, $500$ warm-up steps, $4$ parallel pseudo-documents per batch and $32$ gradient accumulation steps. An early stopping criterion with a patience of $5$ epochs is also applied, according to the d-BLEU evaluated on the validation set. For inference on test sets, the beam size is set to $5$ and the batch size is set to $4$.
[…]
A.5 Detailed Evaluation Results
The paired comparison and the complete BLEU and COMET scores for each test set are given in Tables 11, 12, 13, 14 and 15.
Document-level Fine-tuning
Table 11 reports average differences of automatic scores between fine-tuned MT systems and the corresponding pretrained models (NLLB or TowerBase), for varying test document lengths. ds-BLEUs are averaged over $52$ complete TED talks and COMET scores are averaged over $5,103$ sentences. Fine-tuning significantly improves over base conditions for all lengths, with larger increases for longer test texts, where the baseline scores were initially very poor. Both metrics yield consistent conclusions, except for the sentence-level assessment of NLLB fine-tuned on TED-U, which is slightly worse than the baseline according to ds-BLEU (-$0.8$), but for which COMET detects no difference. For TowerBase, DLFT is always beneficial.
ds-BLEU COMET
FT-U FT-G FT-U FT-G
NLLB sent -0.8 (0.00) -0.8 (0.01) - -0.3 (0.01)
256 7.6 (0.00) 8.0 (0.00) 13.0 (0.00) 13.8 (0.00)
512 26.3 (0.00) 27.0 (0.00) 33.4 (0.00) 33.8 (0.00)
768 33.5 (0.00) 28.0 (0.00) 39.0 (0.00) 27.8 (0.00)
1024 33.5 (0.00) 21.2 (0.00) 41.9 (0.00) 22.6 (0.00)
TowerBase sent 2.4 (0.00) 2.6 (0.00) 0.7 (0.00) 0.7 (0.00)
256 2.1 (0.00) 1.6 (0.00) 2.3 (0.00) 2.3 (0.00)
512 2.1 (0.00) 2.0 (0.01) 2.4 (0.00) 2.6 (0.00)
768 3.5 (0.00) 3.2 (0.00) 3.7 (0.00) 3.7 (0.00)
1024 5.6 (0.00) 5.0 (0.00) 5.6 (0.00) 5.5 (0.00)
1200 5.4 (0.00) 5.1 (0.00) 5.5 (0.00) 5.4 (0.00)
1600 8.4 (0.00) 8.3 (0.00) 10.1 (0.00) 10.3 (0.00)
2048 6.8 (0.00) 6.1 (0.00) 8.8 (0.00) 7.9 (0.00)
Table 11: Average difference (and p-values) in ds-BLEU or 100×100$\times$100 ×COMET between fine-tuned models (FT) and the corresponding pretrained models NLLB (top) and TowerBase (bottom). U and G denote the corpora TED-U and TED-G respectively. - for p-values $>0.05$. Positive values indicate that the fine-tuned model improves over the baseline.
2014 2015 2016 2017
sent FT 43.8 (0.98) 44.0 (0.99) 40.4 (1.00) 41.3 (1.00)
Unif 42.4 (0.98) 42.5 (0.99) 39.7 (1.00) 40.2 (1.00)
SHAPE 40.7 (0.97) 41.5 (0.98) 38.3 (1.00) 39.4 (1.00)
256 FT 43.5 (0.98) 43.1 (0.99) 40.4 (1.00) 40.8 (1.00)
Unif 43.9 (0.98) 43.2 (0.99) 39.7 (1.00) 40.6 (1.00)
SHAPE 43.3 (0.98) 43.2 (0.99) 39.8 (1.00) 40.2 (0.99)
512 FT 43.5 (0.98) 43.4 (0.98) 40.2 (1.00) 40.4 (1.00)
Unif 43.8 (0.98) 43.8 (0.99) 39.8 (1.00) 41.0 (1.00)
SHAPE 40.6 (0.91) 40.5 (0.92) 37.0 (0.9) 40.4 (0.98)
768 FT 36.6 (0.87) 36.4 (0.88) 35.3 (0.92) 35.6 (0.93)
Unif 41.8 (0.95) 39.1 (0.88) 38.1 (0.95) 39.4 (0.97)
SHAPE 34.0 (0.75) 34.9 (0.77) 33.8 (0.84) 34.9 (0.85)
1024 FT 28.6 (0.70) 29.1 (0.75) 28.9 (0.80) 28.7 (0.79)
Unif 36.1 (0.81) 38.7 (0.87) 34.9 (0.87) 37.2 (0.92)
SHAPE 32.4 (0.73) 30.9 (0.69) 30.8 (0.75) 28.0 (0.69)
1200 FT 25.2 (0.64) 25.8 (0.74) 24.1 (0.71) 24.3 (0.73)
Unif 34.6 (0.80) 36.4 (0.82) 30.0 (0.74) 32.4 (0.80)
SHAPE 27.2 (0.61) 30.3 (0.68) 24.6 (0.64) 23.9 (0.60)
1600 FT 18.2 (0.53) 19.3 (0.62) 19.2 (0.62) 19.6 (0.59)
Unif 25.5 (0.59) 30.1 (0.70) 26.9 (0.68) 29.4 (0.72)
SHAPE 22.0 (0.50) 21.7 (0.50) 22.1 (0.56) 20.6 (0.57)
2048 FT 15.4 (0.49) 12.4 (0.52) 16.7 (0.58) 14.8 (0.61)
Unif 22.0 (0.51) 21.6 (0.50) 24.3 (0.60) 20.6 (0.53)
SHAPE 18.7 (0.43) 15.1 (0.44) 14.6 (0.38) 13.4 (0.48)
2014 2015 2016 2017
sent FT 84 85 85 84
Unif 82 84 84 83
SHAPE 82 83 83 83
256 FT 81 82 82 81
Unif 81 82 82 81
SHAPE 80 82 81 80
512 FT 81 82 81 80
Unif 81 82 81 81
SHAPE 77 78 76 77
768 FT 69 69 70 69
Unif 74 75 74 73
SHAPE 68 69 68 68
1024 FT 58 59 60 58
Unif 67 65 67 67
SHAPE 60 62 61 59
1200 FT 56 56 55 53
Unif 61 65 62 60
SHAPE 55 59 55 54
1600 FT 50 49 49 49
Unif 55 58 54 56
SHAPE 51 52 50 48
2048 FT 47 42 48 45
Unif 51 49 51 48
SHAPE 46 43 47 44
Table 12: ds-BLEU (and brevity penalty) (left) and 100×100$\times$100 ×COMET (right) scores for FT-NLLB-G (FT), Unif-NLLB-G (Unif), and SHAPE-NLLB-U (SHAPE) trained on TED-G with target max source document length $M=2048$.
2014 2015 2016 2017
sent FT 44.2 (0.99) 43.5 (0.99) 40.4 (1.00) 41.4 (1.00)
Unif 40.1 (0.95) 40.4 (0.97) 38.0 (1.00) 38.1 (0.99)
SHAPE 39.4 (0.97) 40.0 (0.98) 36.3 (1.00) 37.8 (0.99)
256 FT 43.2 (0.98) 42.8 (0.99) 39.7 (1.00) 40.5 (1.00)
Unif 42.8 (0.98) 42.4 (0.99) 39.5 (1.00) 40.3 (1.00)
SHAPE 42.0 (0.97) 42.5 (0.99) 39.6 (1.00) 39.4 (1.00)
512 FT 42.9 (0.98) 43.1 (0.99) 39.2 (1.00) 39.4 (1.00)
Unif 43.4 (0.98) 43.0 (0.99) 39.8 (1.00) 40.5 (1.00)
SHAPE 39.7 (0.89) 41.1 (0.94) 38.2 (0.95) 39.0 (0.98)
768 FT 43.5 (0.98) 43.0 (0.99) 39.4 (1.00) 39.8 (1.00)
Unif 44.0 (0.98) 43.7 (0.99) 40.4 (1.00) 40.9 (1.00)
SHAPE 39.6 (0.88) 41.4 (0.93) 37.4 (0.93) 36.8 (0.91)
1024 FT 42.6 (0.96) 42.6 (0.97) 39.2 (1.00) 39.6 (1.00)
Unif 42.6 (0.96) 44.1 (0.98) 39.6 (1.00) 40.6 (0.99)
SHAPE 40.3 (0.91) 42.8 (0.96) 36.8 (0.92) 37.8 (0.94)
1200 FT 38.3 (0.88) 39.3 (0.91) 36.3 (0.92) 36.4 (0.92)
Unif 39.5 (0.89) 43.0 (0.97) 38.0 (0.95) 39.2 (0.98)
SHAPE 36.9 (0.84) 37.2 (0.87) 32.6 (0.82) 34.3 (0.86)
1600 FT 31.5 (0.77) 30.4 (0.73) 30.9 (0.83) 30.3 (0.80)
Unif 31.5 (0.72) 34.4 (0.82) 34.2 (0.88) 33.8 (0.84)
SHAPE 28.0 (0.68) 29.4 (0.68) 31.1 (0.79) 31.7 (0.82)
2048 FT 27.4 (0.69) 24.0 (0.63) 26.7 (0.75) 23.6 (0.71)
Unif 30.2 (0.68) 25.3 (0.60) 31.5 (0.79) 28.0 (0.68)
SHAPE 26.1 (0.63) 27.9 (0.66) 26.8 (0.69) 26.9 (0.74)
2014 2015 2016 2017
sent FT 84 85 85 84
Unif 82 83 83 83
SHAPE 81 82 82 82
256 FT 80 81 81 81
Unif 80 81 81 81
SHAPE 79 81 81 80
512 FT 80 81 81 81
Unif 81 82 81 81
SHAPE 75 79 79 78
768 FT 80 81 81 81
Unif 81 82 82 81
SHAPE 75 78 75 75
1024 FT 78 79 78 78
Unif 80 81 81 80
SHAPE 74 78 75 73
1200 FT 71 73 70 69
Unif 76 79 76 75
SHAPE 68 70 69 66
1600 FT 61 60 61 60
Unif 62 61 61 64
SHAPE 57 59 62 59
2048 FT 57 53 57 54
Unif 59 57 57 54
SHAPE 51 52 57 53
Table 13: ds-BLEU (and brevity penalty) (left) and 100×100$\times$100 ×COMET (right) scores for FT-NLLB-U (FT) Unif-NLLB-U (Unif), and SHAPE-NLLB-U (SHAPE) trained on TED-U with target max source document length $M=2048$.
2014 2015 2016 2017
sent FT 46.5 (0.98) 45.1 (0.99) 42.3 (1.00) 41.0 (1.00)
Unif 46.5 (0.98) 45.0 (0.99) 42.3 (1.00) 41.1 (1.00)
SHAPE 46.4 (0.98) 45.2 (0.99) 42.4 (1.00) 41.2 (1.00)
256 FT 44.6 (0.98) 45.1 (0.99) 42.3 (1.00) 41.9 (1.00)
Unif 44.5 (0.99) 45.1 (0.99) 42.2 (1.00) 41.8 (1.00)
SHAPE 46.2 (0.98) 45.2 (0.99) 42.4 (1.00) 42.1 (1.00)
512 FT 43.7 (0.98) 45.0 (1.00) 41.4 (1.00) 41.6 (1.00)
Unif 43.7 (0.98) 44.8 (1.00) 41.4 (1.00) 41.4 (1.00)
SHAPE 45.5 (0.98) 45.1 (0.99) 41.6 (1.00) 41.6 (1.00)
768 FT 44.0 (0.98) 44.2 (0.99) 40.3 (1.00) 40.7 (1.00)
Unif 44.0 (0.98) 44.2 (0.99) 40.2 (1.00) 40.7 (1.00)
SHAPE 45.6 (0.98) 44.4 (0.99) 41.3 (1.00) 41.5 (1.00)
1024 FT 42.7 (0.98) 40.6 (0.99) 38.9 (1.00) 40.4 (1.00)
Unif 42.2 (0.97) 42.6 (0.99) 39.4 (1.00) 40.1 (1.00)
SHAPE 44.5 (0.98) 40.9 (0.99) 39.2 (1.00) 39.7 (1.00)
1200 FT 42.6 (0.98) 43.0 (0.99) 38.9 (1.00) 40.8 (1.00)
Unif 42.5 (0.98) 42.7 (0.99) 39.3 (1.00) 40.6 (1.00)
SHAPE 44.0 (0.98) 42.8 (0.99) 39.3 (1.00) 40.6 (1.00)
1600 FT 42.0 (0.97) 41.1 (0.98) 37.0 (1.00) 38.7 (1.00)
Unif 42.0 (0.97) 40.0 (0.98) 37.5 (1.00) 37.3 (1.00)
SHAPE 42.2 (0.97) 40.6 (0.98) 38.5 (1.00) 39.2 (0.99)
2048 FT 33.0 (0.99) 32.5 (0.99) 31.6 (1.00) 29.2 (0.97)
Unif 33.1 (0.99) 33.7 (0.99) 32.2 (0.98) 26.5 (0.97)
SHAPE 34.1 (0.97) 35.1 (0.98) 32.1 (1.00) 30.2 (0.97)
2014 2015 2016 2017
sent FT 85 86 85 85
Unif 85 86 85 85
SHAPE 85 86 85 85
256 FT 83 84 83 83
Unif 83 84 83 83
SHAPE 83 84 83 83
512 FT 82 84 83 82
Unif 82 84 83 82
SHAPE 82 84 83 82
768 FT 82 83 83 82
Unif 82 83 83 82
SHAPE 82 84 83 82
1024 FT 81 81 82 81
Unif 81 82 83 81
SHAPE 81 81 83 81
1200 FT 78 82 81 81
Unif 78 82 81 81
SHAPE 80 82 82 81
1600 FT 80 79 79 80
Unif 79 78 79 78
SHAPE 79 79 80 79
2048 FT 68 69 68 64
Unif 69 70 70 62
SHAPE 65 73 70 68
Table 14: ds-BLEU (and brevity penalty) (left) and 100×100$\times$100 ×COMET (right) scores for FT-Tower-G (FT), Unif-Tower-G (Unif) and SHAPE-Tower-G (SHAPE) trained on TED-G with target max source document length 2048 ($M=4096$).
2014 2015 2016 2017
sent FT 46.2 (0.99) 45.1 (0.99) 42.1 (1.00) 40.8 (1.00)
Unif 46.4 (0.98) 45.1 (0.99) 42.2 (1.00) 40.9 (1.00)
SHAPE 46.3 (0.98) 45.2 (0.99) 42.4 (1.00) 41.0 (1.00)
256 FT 46.3 (0.98) 45.1 (0.99) 42.3 (1.00) 41.8 (1.00)
Unif 44.5 (0.98) 45.2 (0.99) 42.3 (1.00) 41.9 (1.00)
SHAPE 46.0 (0.98) 45.3 (0.99) 42.4 (1.00) 42.0 (1.00)
512 FT 44.4 (0.98) 44.9 (0.99) 41.2 (1.00) 41.6 (1.00)
Unif 43.0 (0.98) 45.0 (0.99) 41.4 (1.00) 41.6 (1.00)
SHAPE 44.5 (0.98) 45.0 (0.99) 41.3 (1.00) 41.6 (1.00)
768 FT 44.1 (0.98) 44.6 (0.99) 41.1 (1.00) 40.8 (1.00)
Unif 43.2 (0.99) 44.5 (0.99) 41.4 (1.00) 41.0 (1.00)
SHAPE 44.2 (0.99) 44.5 (0.99) 41.2 (1.00) 41.7 (1.00)
1024 FT 43.9 (0.97) 41.6 (0.99) 39.6 (1.00) 39.7 (1.00)
Unif 42.7 (0.98) 43.8 (0.99) 39.7 (1.00) 39.6 (1.00)
SHAPE 44.1 (0.98) 42.9 (0.99) 39.8 (1.00) 39.5 (1.00)
1200 FT 43.2 (0.97) 42.9 (0.99) 39.5 (1.00) 40.8 (1.00)
Unif 43.2 (0.98) 43.1 (0.99) 38.2 (1.00) 40.9 (1.00)
SHAPE 43.8 (0.98) 43.2 (0.99) 39.2 (1.00) 39.7 (1.00)
1600 FT 41.6 (0.95) 40.5 (0.97) 37.6 (1.00) 39.8 (0.99)
Unif 41.1 (0.97) 41.6 (0.98) 36.5 (1.00) 38.0 (1.00)
SHAPE 42.6 (0.96) 42.1 (0.98) 37.3 (1.00) 40.5 (1.00)
2048 FT 34.4 (0.96) 35.3 (0.97) 31.2 (0.99) 28.2 (0.96)
Unif 34.6 (0.97) 35.5 (0.98) 30.2 (1.00) 30.9 (0.97)
SHAPE 35.2 (0.97) 34.6 (0.95) 31.9 (0.99) 31.5 (0.96)
2014 2015 2016 2017
sent FT 85 86 85 85
Unif 85 86 85 85
SHAPE 85 86 85 86
256 FT 83 84 83 83
Unif 83 84 83 82
SHAPE 83 84 83 82
512 FT 82 84 83 82
Unif 82 84 83 82
SHAPE 82 84 83 82
768 FT 82 83 83 82
Unif 82 84 83 82
SHAPE 82 83 83 82
1024 FT 81 81 83 81
Unif 81 83 83 81
SHAPE 81 82 83 81
1200 FT 80 82 81 81
Unif 80 82 81 81
SHAPE 80 82 81 80
1600 FT 79 79 80 80
Unif 80 80 78 79
SHAPE 79 81 78 80
2048 FT 68 72 69 64
Unif 70 72 69 67
SHAPE 70 73 70 68
Table 15: ds-BLEU (and brevity penalty) (left) and 100×100$\times$100 ×COMET (right) for FT-Tower-U (FT), Unif-Tower-U (Unif) and SHAPE-Tower-U (SHAPE) trained on TED-U with target max source document length 2048 ($M=4096$).
TED-U TED-G FT Unif SHAPE
FT vs Unif FT vs SHAPE FT vs Unif FT vs SHAPE U vs G U vs G U vs G
sent -0.1 (0.20) -0.2 (0.01) -0.0 (0.60) -0.1 (0.18) -0.1 (0.05) -0.1 (0.19) -0.1 (0.36)
256 0.5 (0.32) -0.0 (0.87) 0.1 (0.22) -0.5 (0.24) 0.5 (0.32) 0.1 (0.21) -0.1 (0.48)
512 0.3 (0.55) -0.1 (0.10) 0.1 (0.08) -0.6 (0.23) 0.1 (0.87) -0.1 (0.82) -0.4 (0.28)
768 0.2 (0.46) -0.2 (0.67) 0.0 (0.84) -0.9 (0.05) 0.3 (0.11) 0.2 (0.65) -0.4 (0.19)
1024 -0.2 (0.82) -0.3 (0.49) -0.4 (0.44) -0.5 (0.35) 0.6 (0.29) 0.4 (0.34) 0.5 (0.38)
1200 0.2 (0.44) 0.1 (0.82) 0.0 (0.70) -0.4 (0.06) 0.3 (0.24) 0.1 (0.71) -0.2 (0.50)
1600 0.6 (0.45) -0.7 (0.01) 0.4 (0.58) -0.5 (0.40) 0.1 (0.84) 0.0 (1.00) 0.4 (0.49)
2048 -0.5 (0.61) -1.0 (0.22) 0.2 (0.84) -1.3 (0.19) 0.7 (0.46) 1.3 (0.13) 0.4 (0.44)
sent -0.0 (0.52) -0.0 (0.93) 0.0 (0.38) 0.0 (0.60) -0.0 (0.95) 0.0 (0.24) 0.0 (0.56)
256 0.0 (0.60) 0.0 (0.85) 0.0 (0.73) -0.1 (0.07) -0.0 (0.63) -0.1 (0.46) -0.2 (0.05)
512 -0.1 (0.15) -0.2 (0.04) 0.1 (0.16) -0.1 (0.27) -0.2 (0.09) 0.0 (0.76) -0.0 (0.62)
768 -0.1 (0.29) 0.1 (0.28) -0.1 (0.34) -0.3 (0.00) 0.0 (0.87) 0.0 (0.69) -0.4 (0.00)
1024 -0.6 (0.22) -0.5 (0.32) -0.3 (0.46) -0.2 (0.09) 0.1 (0.45) 0.3 (0.10) 0.4 (0.42)
1200 0.1 (0.69) 0.2 (0.51) 0.1 (0.50) -0.4 (0.16) 0.1 (0.61) 0.0 (0.93) -0.4 (0.12)
1600 0.2 (0.69) -0.4 (0.27) 0.6 (0.37) 0.1 (0.70) -0.2 (0.62) 0.2 (0.73) 0.4 (0.49)
2048 -0.9 (0.40) -1.3 (0.16) -0.2 (0.84) -1.7 (0.13) 0.9 (0.50) 1.5 (0.10) 0.4 (0.47)
Table 16: Average difference (and p-values) in ds-BLEU (top) evaluated on full TED talks and 100×100$\times$100 ×COMET (bottom) evaluated on realigned sentences for TowerBase. Left and middle: paired comparison between the original fine-tuning (FT), unifPE and SHAPE on TED-U and TED-G respectively. Right: differences between fine-tuning on TED-U (U) and TED-G (G). A positive value indicates that in the comparison pair, the translation of the first item achieves higher scores than that of the second. Significant differences with p-values $<0.05$ are in bold.
Realignment Issues
Since the COMET score is sentence-based, its computation requires a realignment between hypotheses and reference sentences in the Doc2Doc scenario. However, due to issues with long documents, translation hypotheses can be incomplete, resulting in empty alignment for some sentences. These untranslated sentences often occur in the final part of long documents. Table 10 displays the statistics of empty alignments across all the $5,103$ sentences. This issue is more severe for NLLB models than TowerBase models, which is consistent with the poor BP reported in Tables 12 and 13.
disabledisabletodo: disablechange all BLEU score to ds-BLEU at corpus leveldisabledisabletodo: disablecheck again the remarks of Paul and Rachel